I need a step by step with all the details, I need all the parts and you would be saving my butt, thank you in advanced.

2. Three resistors, RI-, Rs 15 2 are 20 a Assume that AV = IS V, and RRs - 30 o connected in a circuit as shown below. through each resistor. (C) Determine the voltage across each resistor. (A) Calculate the equivalent resistance of the circuit. (B) find the current WWW AV
Answer #1

a)

R2 and R3 are in parallel

so, R23=R2xR3/(R2+R3)

=15x30/(15+30)

=10 ohm

Now R23 and R1 are in series, So

Reff= R23+R1

=10+20

= 30 ohm

ans- Equivalent resistance of circuit is 30 ohm

b)

Current I in the circuit will be. I=V/Reff

=15/30

I=0.5 A

Since R1 is in series with R2 and R3 resistor

So, I1= 0.5 A

Now, I2=0.5/2

=0.25A

Similarly , I3= 0.25A

Ans- current through R1, I1=0.5 A

Current through R2, I2=0.25A

Current through R3, I3= 0.25 A

c)

In parallel, voltage is same and in series voltage is divided.

V1= IxR1

=0.5x20

=10 v

V2= I2xR2

=0.25x15

V2=3.75 V

V3=I3xR3

= 0.25x30

=7.50

Ans- Voltage through R1, V1=10 V

V2= 3.75v V

V3= 7.5 V